What is the molarity of H3PO4 solution if 358 mL is completely titrated by 876 mL of 0.0102 M Ba(OH)2 solution?



1
Expert's answer
2021-06-02T06:30:20-0400

"H_\n3\nP\nO_\n4\n(\na\nq\n)\n+\nB\na\n(\nO\nH\n)_\n2\n(\na\nq\n)\n\u2192\nB\na\nH\nP\nO_\n4\n(\na\nq\n)\n+\n2\nH_\n2\nO\n(\nl\n)"


Now moles of "Ba(OH)_2=0.0102 \times876 \times10^{-3} = 8.94 \times 10^{-3} moles"


As the ratio of "Ba(OH)_2" and "H_3PO_4" is 1:1 so the moles of "H_3PO_4=8.94 \times 10^{-3}"


Now Molarity of "H_3PO_4 =" "\dfrac{moles}{volume(ml )}\times1000"


"=" "\dfrac{8.94\times10^{-3}}{358} \times 1000=0.025M"